A 60 feet long piece of string is cut into 3 pieces. The second piece is twice as long as the first piece. The third piece is 5 feet longer than the second piece. What is the length of the shortest piece of rope?
Answer by hinaahsan(5)

let the first piece be x of length
let the second piece be 2x of length
let the third piece be 2x+5 of length
total length = 60
x+2x+2x+5=60
5x+5=60
5x=60-5
5x=55
x=55/5
x=11
length of 1st piece = 11
length of 2nd piece =22
length of 3rd piece =27
counter check
11+22+27=60